The volume occupied by 1.008 gm of Hydrogen at STP is a. 22.4cc b. 11.2cc c. 22.4l d. 11.2l

Grade:

2 Answers

Arun
4 years ago
Now, as one mole of any gas occupies 22.4 litres of volume at STP, 1 mole of Hydrogen occupies 22.4 litres. Hence, 2 gm of Hydrogen occupies 22.4 litres of volume. So it is pretty obvious that 1 gm of Hydrogen gas occupies 11.2 litres at STP
Puja Nepal
3 years ago
1 gm = 1000 mg
So, 1.008 mg = 1.008/1000
 
Weight in gm/ molar mass = volume/  22.4 l
(1.008/1000) / 2 = V/ 22.4 l
 
 
So, v = 11.2 l